OpenStudy (anonymous):

Choose the correct simplification of the expression (5x4y2z2)(3x4y3z5). 15x16y6z10 15x8y5z7 8x8y5z7 8x16y6z

OpenStudy (johnweldon1993):

Lets see what we got here :P \[\large (5x^4 y^2 z^2)(3x^4 y^3 z^5)\] right?

OpenStudy (anonymous):

yess sirr. c:

OpenStudy (johnweldon1993):

Alright...so all we need to do is multiply the terms together we can eliminate 2 options off the bat because we know 5 times 3 = 15 right so C and D are gone...

OpenStudy (johnweldon1993):

Nowwww, when we multiply exponents with the same base example x^4 and x^4 here...what do we do? or what do we get?

OpenStudy (anonymous):

16?

OpenStudy (anonymous):

is that what we get or no?

OpenStudy (johnweldon1993):

Not quite, the rules of exponents actually say that \[\large x^4\times x^4 = x^{4 + 4}\]

OpenStudy (johnweldon1993):

We only multiply the exponents if we have something like \(\large (x^4)^4\) when they are directly separated by a parenthesis...then we multiply but otherwise *and here* we add the exponents... so \[\large x^4 \times x^4 = x^{4 + 4} =x^8\]

OpenStudy (anonymous):

oh.

OpenStudy (johnweldon1993):

And we can keep going *theres only 1 answer choice with x^8 on here but for practice lol* \[\large y^2 \times y^3 = ?\] and \[\large z^2 \times z^5 = ?\]

OpenStudy (anonymous):

y=6 ?

OpenStudy (anonymous):

idk this all confuses me

OpenStudy (johnweldon1993):

Not quite...we need to add the exponents So imagine we are doing something easier like \[\large (2^6)(2^3)\] and we want to know what that is...now we see that the base number (2) is the same in both cases...so we are going to leave it alone...and we are going to just add the exponents so \[\large (2^6)(2^3) \rightarrow 2^{6 + 3} \rightarrow 2^9\] and it works the same here

OpenStudy (johnweldon1993):

So here...we just focus on say the 'y' terms we have \[\large (y^2)(y^3)\] this time...the base number *which is 'y'* is the same in both cases...so we leave it alone...and just add the exponents \[\large (y^2)(y^3) \rightarrow y^{2 + 3} \rightarrow y^5\]

OpenStudy (johnweldon1993):

Is that making a little more sense maybe?

OpenStudy (anonymous):

okay. yeah i get it little bit more now.

OpenStudy (johnweldon1993):

Okay, so test time :P If we have \[\large (z^2)(z^5)\] what would that be?

OpenStudy (anonymous):

\[z ^{7}\]

OpenStudy (johnweldon1993):

Perfect :) Alright good... so now it is confirmed that our simplification was \[\large 15x^8y^5z^7\]

OpenStudy (anonymous):

thank you.

OpenStudy (johnweldon1993):

Anytime! If you need any more clarification just tag me back or message me :)